# Mind’s Mirror: Distilling Self-Evaluation Capability and Comprehensive Thinking from Large Language Models

### v2

@@ Abstract
Weize Liu1, Guocong Li1, Kai Zhang2, Bang Du1, Qiyuan Chen1 Xuming Hu3††thanks: Corresponding authors., Hongxia Xu2,411footnotemark: 1, Jintai Chen5, Jian Wu2,4 1Zhejiang University 2School of Public Health, Zhejiang University 3The Hong Kong University of Science and Technology (Guangzhou) 4Liangzhu Laboratory and Institute of Wenzhou, Zhejiang University 5Computer Science Department, University of Illinois Urbana-Champaign {weizeliu1115,xuminghu97}@gmail.com einstein@zju.edu.cn
@@ 4.3 Effect of the number of CoTs
Model Input | Premise: East Bengal Football Club is a professional football club, based in Kolkata, West Bengal, India. It currently competes in the I-League, the top tier of Indian football. The club has won three National Football League (India) (later named as I league) titles, eight Federation Cups, and three Indian Super Cups, among others. The club is the current holder of the Calcutta Football League. |
Hypothesis: All of the team members live in West Bengal. | |
@@ Appendix D Self-evaluation prediction accuracy and evaluation outputs of SLMs
Model Input | Every day Ryan spends 6 hours on learning English 7 hours on learning Chinese and 3 hours on learning Spanish. How many hours does he spend on learning English, Chinese, and Spanish in all? |
@@
Model Input | Question: As a result of dying, what happens to organic material? |
Choices: [change of color, stop breathing, wake up, death and decay, getting cold] | |
@@
Label: change of color | |
@@
Model Input | Paige and her mom planted 10 seeds in each flower bed, and there are 45 flower beds. So, they planted a total of 10 * 45 = 450 seeds. The answer is (10 * 45). |

### v3

@@ Abstract
Weize Liu, Guocong Li, Kai Zhang, Bang Du, Qiyuan Chen, Xuming Hu††thanks: Corresponding authors., Hongxia Xu11footnotemark: 1, Jintai Chen, Jian Wu Zhejiang University School of Public Health, Zhejiang University The Hong Kong University of Science and Technology (Guangzhou) Liangzhu Laboratory and Institute of Wenzhou, Zhejiang University Computer Science Department, University of Illinois Urbana-Champaign {weizeliu1115,xuminghu97}@gmail.com einstein@zju.edu.cn
@@ 4.3 Effect of the number of CoTs
Model Input |
Premise: East Bengal Football Club is a professional football club, based in Kolkata, West Bengal, India. It currently competes in the I-League, the top tier of Indian football. The club has won three National Football League (India) (later named as I league) titles, eight Federation Cups, and three Indian Super Cups, among others. The club is the current holder of the Calcutta Football League. |
Hypothesis: All of the team members live in West Bengal. |
@@ Appendix D Self-evaluation prediction accuracy and evaluation outputs of SLMs
Model Input |
Every day Ryan spends 6 hours on learning English 7 hours on learning Chinese and 3 hours on learning Spanish. How many hours does he spend on learning English, Chinese, and Spanish in all? |
@@
Model Input |
Question: As a result of dying, what happens to organic material? |
Choices: [change of color, stop breathing, wake up, death and decay, getting cold] |
@@
Label: change of color |
@@
Model Input |
Paige and her mom planted 10 seeds in each flower bed, and there are 45 flower beds. So, they planted a total of 10 * 45 = 450 seeds. The answer is (10 * 45). |